# Audit Cards: Contextualizing AI Evaluations

### v1

@@ 1 Introduction
AI governance frameworks [^1] This includes frontier AI safety policies from AI companies (e.g. OpenAI Preparedness Framework), regulations (e.g. EU AI Act), and other norms (e.g. NIST AI Risk Management Framework). See Appendix D for a complete list. are being designed to increasingly rely on audits. [^2] We understand audits as formalized evaluation processes of AI models. Usually, a specific evaluation procedure (e.g. benchmarking, red-teaming, uplift studies) underlies the audit. We refer to these underlying procedures as evaluations. See Appendix A for details. Within these frameworks, audits are meant to (1) identify potential risks, (2) incentivize more responsive development practices, and (3) involve more stakeholders in the system deployment process. However, audits can only fulfill this role effectively if the underlying evaluations are both technically rigorous and effectively integrated into decision-making (Hardy et al., 2024). However, not all evaluations are equally rigorous (e.g., Raji et al., 2022; Birhane et al., 2024; Mökander, 2023; Anderljung et al., 2023; Kolt et al., 2024; Casper et al., 2024; Reuel et al., 2024b; Li & Goel, 2024).
To date, much prior literature has focused on the technical side of audits (see Section 2). However, evaluations are never conducted in a vacuum (Wallach et al., 2025). Their results are intimately shaped, not only by the technical methods employed, but by numerous non-technical details of their context and design (see Section 3). For example, while the research field has long understood the importance of conflict of interest disclosures for academic integrity (e.g., Knerr & D’Amelia, 2020), auditor independence remains a consistent concern in the AI ecosystem (Costanza-Chock et al., 2022; Evans et al., 2023).
Even technically rigorous evaluations can be uninformative or even actively misleading if reported on selectively or obscurely (Ananny & Crawford, 2018). This risk is magnified given the vested interest of AI developers in obtaining favorable audit results. In AI and other fields alike, companies undergoing evaluations have strong incentives to game audits and influence reporting in potentially misleading ways (e.g., Krawiec, 2003; Lu, 2006; Marquis et al., 2016).
@@
Figure 1: Our audit card template requires reporting on three principles and six features. (Left) The three principles (justification, limitations, and assumptions) further transparency about the process behind an audit. (Right) The six features offer key methodological and contextual information. See Section 3 and Section C.1 for details.
Thus, for AI audits to play a meaningful role in governance, evaluation reporting must be done in a way that minimizes the chance of omitting key contextual information. To address this challenge, we propose “audit cards”. Akin to model and system cards (e.g., Mitchell et al., 2019), audit cards provide a structured format for reporting key background, design choices, and justifications. They are designed to offer a standardized approach to reporting that enhances transparency and facilitates informed interpretations of results. This need not necessarily be reported in a separate audit card document, rather the relevant context can be added to existing reports, such as sections in model and system cards.
@@
    Audit cards and survey of evaluation literature: Based on a survey of 25 prior works on AI evaluations (Section 3), we propose a template and checklist (Appendix C) for AI audit cards including three key principles (justification, limitations, and assumptions) and six key types of information (auditor identity, evaluation scope, methodology, resource access, process integrity, and review mechanisms).
@@
    Survey of frontier AI evaluation reports: We analyze the thoroughness of 20 existing evaluation reports for frontier systems (Section 4).
@@ 2 Related work
Transparency and reporting: In the past decade, frontier AI research and development have largely shifted from being predominantly academia-driven to predominantly industry-driven (Maslej et al., 2024). This shift has been accompanied by a decrease in transparency around proprietary state-of-the-art systems. As such, increasing transparency and awareness has emerged as a key goal of AI governance (e.g., Felzmann et al., 2020; Haresamudram et al., 2023; Winecoff & Bogen, 2024; Chan et al., 2024; Bommasani et al., 2024; Kolt et al., 2024). One barrier to meaningful transparency is a lack of standards for reporting (Maslej et al., 2024). In response to this challenge, previous works have proposed data, model, system, agent, and usage “cards” (Pushkarna et al., 2022; Mitchell et al., 2019; Gursoy & Kakadiaris, 2022; Casper et al., 2025; Wahle et al., 2023) for documenting key information within the AI ecosystem. In this paper, we build on past work by introducing the notion of an “audit card,” surveying what prior literature suggests they should contain, and analyzing the current state of reporting around audits.
@@
AI audits: Formal evaluations, also known as “audits,” of AI systems have been proposed as a key objective to facilitate transparency and scrutiny (Raji et al., 2022; Anderljung et al., 2023; Mökander, 2023; Li & Goel, 2024; Reuel et al., 2024a). Meanwhile, audits are increasingly incorporated into frameworks for AI governance (e.g., EU AI Act, 2024). However, as we will show in Section 5, current governance frameworks often lack substantial guidance for how to report on audits.
@@
Technically rigorous evaluations: AI systems are audited using a variety of approaches including case studies, benchmarks, red-teaming, and mechanistic analysis (Bengio et al., 2024). However, the science of evaluating AI systems is still nascent (Apollo Research, 2024). Not all evaluations are equally technically rigorous, and their apparent outcomes can be highly sensitive to framing and design (e.g., Schaeffer et al., 2023; Burnell et al., 2023; Khan et al., 2025). Toward improved technical practices, Reuel et al. (2024b) outline a set of 46 best practices for rigorous AI benchmark design. Meanwhile, METR (2024b) work toward a portable standard for framing and conducting capability evaluations. However, unlike prior work, here we focus on rigor in reporting both key technical details and context behind AI audits.
@@
On the inherent sociotechnical nature of evaluations: AI evaluations depend on technical tools to assess system properties and risks. However, they are never conducted in a vacuum; they are always embedded in a broader sociotechnical and political context. Meanwhile, the standards that systems are evaluated against are inherently based on subjective human values. Thus, AI evaluations represent a social science measurement challenge (Wallach et al., 2025). Aside from lacking technical soundness, evaluations can fail to be rigorous and serve the public’s interest for a variety of nontechnical reasons. Prior works have emphasized the role of evaluation integrity and integration in ensuring meaningful oversight (Ojewale et al., 2024; Raji et al., 2022; Sharkey et al., 2024). By surveying prior literature on evaluation procedures (Section 3), analyzing current evaluation reports (Section 4), and analyzing evaluation frameworks (Section 5), we make progress toward a more contextual and critical understanding of AI audits.
@@ 3 What information do audit cards need for rigorous reporting?
Methodology: We identify key components of audit cards through an initial literature review of work on AI audits and evaluations, transparency, technical evaluation design, and sociotechnical approaches. [^3] We also refined our audit card template based on structured interviews with expert stakeholders. See Section 6. We selected 25 total papers and manually annotated them, scoring the extent to which the paper recommends that aspect of reporting: is it a central argument of the paper (2), a minor argument of the paper (1), or not mentioned in the paper (0) (see further details in Appendix B). In Table 1, we summarize the perspectives from all 25 papers on the three principles and six features. Next, we expand on these principles and features. See Figure 1 for an overview of an audit card and Section C.2 for an actionable audit card checklist that discusses the relative importance of principles and features for different stakeholders and auditing types.
[…]
@@
Table 1: What contextual details do prior works say are key for audit reporting? We divide reporting into three overarching principles (justifications, assumptions, limitations) and six key features (who, what, how, access, integrity, review). We score each paper’s extent of recommendation of each these as 2 (central argument), 1 (minor argument), and 0 (no mention). Refer to Table 4 for our complete, more granular analysis of papers.
3.1 Overarching principles for transparent reporting
@@
The following three overarching principles behind audit design are key for transparency and methodological clarity. These apply across the entire auditing process and the six specific features. We design audit cards to make it easier to adopt these principles throughout the auditing process. This, in return, enables accurate interpretation, facilitates constructive critique, and builds trust in the evaluation process.
@@
Justifications as explicit arguments to support methodological choices, metric selection, and interpretative frameworks. Explicit reporting on justifications enables process transparency. This justification can be through elements of a rigorous scientific process, such as explaining proxies and experimental coverage (Apollo Research, 2024), or through standardized auditing procedures, such as detailed standards identification (Birhane et al., 2024; Costanza-Chock et al., 2022).
@@
Assumptions to articulate the premises underlying the evaluation design and analysis. Reporting on assumptions clarifies the relationship (or lack thereof) between tests and real-world outcomes, presumptions about what constitutes “good” performance, and details about the threat models considered (Reuel et al., 2024b; Barnett & Thiergart, 2024a; b; Raji et al., 2020).
@@
Limitations to explicitly acknowledge the constraints that affect the validity, reliability, or generalizability of findings. Reporting limitations clarifies constraints in data sampling, potential artifacts in the evaluation process, and boundaries of what auditors can reasonably claim based on the evaluation evidence (Mökander et al., 2024; Barnett & Thiergart, 2024b; Reuel et al., 2024b).
3.2 Specific features for contextual information
@@
Aside from overarching principles, reporting on more specific contextual features of audits helps stakeholders assess trustworthiness. We compile recommendations from past literature into six features for reporting evaluation context in audit cards. For a more granular analysis, see Section C.1, where we further divide these six features into constituent aspects.
@@
  • Who are the auditors: Reporting details of auditor identity facilitates trust and clarifies what potential biases might affect their work. It includes reporting the expertise (relevant domain knowledge, accreditations) (Anderljung et al., 2023; METR, b; Reuel et al., 2024b) and background (track record of conducting evaluations, positionality) of auditors (Costanza-Chock et al., 2022; Gebru et al., 2021). Given worries about the privacy of auditors, information about the background can be anonymized and aggregated.
@@
  • What is evaluated: Reporting on evaluation goals enables clear interpretation of findings. This includes the scope (e.g., model vs. scaffolding) and goal (context of use) (Raji et al., 2020; Birhane et al., 2024; Shevlane et al., 2023), type of evaluation (e.g., capability vs. propensity), as the scope of the evaluation influences what the relevant level of access is (METR, 2024b; Liang et al., 2023), as well as reporting on the evaluation obsolescence criteria. [^4] Examples of obsolescence criteria could include substantial amounts of model fine-tuning, changes in system components, or changes in deployment context. This helps stakeholders understand the shelf life of the findings. It also includes reporting whether the evaluation will be repeated, and if so, at what time or after which conditions are met (Barnett & Thiergart, 2024b; Anderljung et al., 2023).
@@
  • How is it evaluated: Explaining the underlying procedures for the evaluation, such as descriptions of the evaluation setup with justifications of how a capability translates to evaluation goal (Apollo Research, 2024; Raji et al., 2020; Reuel et al., 2024b) and provide interpretation of scores (Dobbe et al., 2021; Selbst et al., 2019). The specific relevant aspects to report for this feature may change depending on the underlying evaluation. We therefore refer to the appropriate literature for the underlying evaluation for a complete list of relevant aspects to report and details on how to report on them, e.g. for benchmarks (Reuel et al., 2024b).
@@
  • Access and resources: Report the constraints under which the evaluation was conducted to allow readers to critically assess the audit’s limitations. This includes, what access auditors have to a system (black/white box, with/without safeguards) and which resources available are available (compute, time, funding) as this can have a large impact on the rigor of evaluations (Barnett & Thiergart, 2024b; Ojewale et al., 2024; Casper et al., 2024).
@@
  • Integrity: Accountability for the auditing process establishes trust in auditors and rules out potential conflicts of interest. This includes reporting transparently about the selection process (how auditors were chosen) and compensation (funding sources, incentive structures). These dynamics can significantly affect the evaluation process and are therefore crucial context (Birhane et al., 2024; Costanza-Chock et al., 2022; Gebru et al., 2021).
@@
  • Review and communication: Reports on if and what independent reviews of audit methods facilitate accountability [^5] We discuss issues of intellectual property and non-disclosure agreements (NDAs) standing in the way of internal review in Section 6 and Section 7. This review can be performed by experts at the organization by cross-checking the final report, insofar as they were not involved in the evaluation beforehand. (Reuel et al., 2024b; Anderljung et al., 2023). This further includes contact information and avenues for corrections, executive summaries (covering the evaluation process and result for non-technical audiences), and clear definitions of key terms (Gebru et al., 2021; Bucknall & Trager, 2023; METR, 2024b; Reuel et al., 2024b).
@@
Some of the above features may be considered confidential in certain contexts, with academic researchers potentially sharing different details than commercial entities (Mitchell et al., 2019). Recommendations about how and with whom audit cards should be shared are beyond the scope of this report. However, redactions offer a simple, well-precedented approach for protecting sensitive information (Boone et al., 2015).

### v2

@@ 1 Introduction
AI governance frameworks [^1] This includes frontier AI safety policies from AI companies (e.g. OpenAI Preparedness Framework), regulations (e.g. US AI Action Plan, EU AI Act), and other norms (e.g. NIST AI Risk Management Framework). See Appendix D for a complete list. are being designed to increasingly rely on audits. [^2] We understand audits as formalized evaluation processes of AI models. We refer to the specific evaluation procedure (e.g. benchmarking, red-teaming, uplift studies) underlying the audit as evaluations. See Appendix A for an overview of key terms. Within these frameworks, audits are meant to (1) identify potential risks, (2) incentivize more responsive development practices, and (3) involve more stakeholders in the system deployment process. However, audits can only fulfill this role effectively if the underlying evaluations are both technically rigorous and effectively integrated into decision-making (Hardy et al. 2024). However, not all evaluations are equally rigorous (e.g., Raji et al. 2022; Birhane et al. 2024; Mökander 2023; Anderljung et al. 2023; Kolt et al. 2024; Casper et al. 2024; Reuel et al. 2024b; Li and Goel 2024).
To date, much prior literature has focused on the technical side of audits (see Section 2). However, evaluations are never conducted in a vacuum (Wallach et al. 2025). Their results are intimately shaped, not only by the technical methods employed, but by numerous non-technical details of their context and design (see Section 3). For example, while the research field has long understood the importance of conflict of interest disclosures for academic integrity (e.g., Knerr and D’Amelia 2020), auditor independence remains a consistent concern in the AI ecosystem (Costanza-Chock, Raji, and Buolamwini 2022; Evans et al. 2023).
Even technically rigorous evaluations can be uninformative or actively misleading if reported on selectively or obscurely (Ananny and Crawford 2018). This risk is magnified given the vested interest of AI developers in obtaining favorable audit results. In AI and other fields, companies undergoing evaluations have strong incentives to game audits and influence reporting in potentially misleading ways (e.g., Krawiec 2003; Lu 2006; Marquis, Toffel, and Zhou 2016).
@@
Figure 1: Our audit card template requires reporting on three principles and six features. (Left) The three cross-cutting principles (justification, limitations, and assumptions) further transparency about the process behind an audit. (Right) The six features offer key methodological and contextual information. See Section 3 and Appendix C for details.
Thus, for AI audits to play a meaningful role in governance, evaluation reporting must be done in a way that minimizes the chance of omitting key contextual information. To address this challenge, we propose “audit cards”: a structured reporting framework to document the context of an AI audit rather than the evaluated system itself (Figure 1). Unlike model and system cards (Mitchell et al. 2019), which document system characteristics and capabilities, audit cards address the distinct transparency requirements of evaluation processes-specifically the contextual factors that shape how evaluations are conducted and reported. They are designed to offer a standardized approach to reporting that enhances transparency and facilitates informed interpretations of results. This need not necessarily be reported in a separate audit card document, rather the relevant context can be added as sections in existing model and system cards.
@@
    Audit cards framework addressing evaluation transparency gaps: Based on a survey of 28 prior works on AI evaluations (Section 3), we synthesise existing literature to propose a template and checklist (Appendix C) for AI audit cards that addresses transparency needs not covered by existing tools, including three cross-cutting principles (justification, limitations, and assumptions) and six key types of contextual information (auditor identity, evaluation scope, methodology, resource access, process integrity, and review mechanisms).
@@
    Survey of frontier AI evaluation reports: We analyze the thoroughness of 24 existing evaluation reports for frontier systems (Section 4).
@@ 2 Related work
Transparency and reporting: In the past decade, frontier AI research and development have largely shifted from being predominantly academia-driven to predominantly industry-driven (Maslej et al. 2024). This shift has been accompanied by a decrease in transparency around proprietary state-of-the-art systems. As such, increasing transparency and awareness has emerged as a key goal of AI governance (e.g., Felzmann et al. 2020; Haresamudram, Larsson, and Heintz 2023; Winecoff and Bogen 2024; Chan et al. 2024; Bommasani et al. 2024; Kolt et al. 2024). One barrier to meaningful transparency is a lack of standards for reporting (Maslej et al. 2024). In response to this challenge, previous works have proposed data, model, system, agent, and usage “cards” (Pushkarna, Zaldivar, and Kjartansson 2022; Mitchell et al. 2019; Gursoy and Kakadiaris 2022; Casper et al. 2025; Wahle et al. 2023) for documenting key information within the AI ecosystem. In this paper, we build on past work by introducing the notion of an “audit card,” surveying what prior literature suggests they should contain, and analyzing the current state of reporting around audits.
@@
AI audits: Formal evaluations, also known as “audits,” of AI systems have been proposed as a key objective to facilitate transparency and scrutiny (Raji et al. 2022; Anderljung et al. 2023; Mökander 2023; Li and Goel 2024; Reuel et al. 2024a). Meanwhile, audits are increasingly incorporated into frameworks for AI governance (e.g., EU AI Act 2024). However, as we will show in Section 5, current governance frameworks often lack substantial guidance for how to report on audits.
@@
Technically rigorous evaluations: AI systems are audited using a variety of approaches including case studies, benchmarks, red-teaming, and mechanistic analysis (Bengio et al. 2024). However, the science of evaluating AI systems is still nascent (Apollo Research 2024). Not all evaluations are equally technically rigorous, and their apparent outcomes can be highly sensitive to framing and design (e.g., Schaeffer, Miranda, and Koyejo 2023; Burnell et al. 2023; Khan, Casper, and Hadfield-Menell 2025). Toward improved technical practices, Reuel et al. (2024b) outline a set of 46 best practices for rigorous AI benchmark design. Meanwhile, METR (2024) work toward a portable standard for framing and conducting capability evaluations. However, unlike prior work, here we focus on rigor in reporting both key technical details and context behind AI audits.
@@
On the inherent sociotechnical nature of evaluations: AI evaluations depend on technical tools to assess system properties and risks. However, they are never conducted in a vacuum; they are always embedded in a broader sociotechnical and political context. Meanwhile, the standards that systems are evaluated against are inherently based on subjective human values. Thus, AI evaluations represent a social science measurement challenge (Wallach et al. 2025). Aside from lacking technical soundness, evaluations can fail to be rigorous and serve the public’s interest for a variety of nontechnical reasons. Prior works have emphasized the role of evaluation integrity and integration in ensuring meaningful oversight (Ojewale et al. 2024; Raji et al. 2022; Sharkey et al. 2024). By surveying prior literature on evaluation procedures (Section 3), analyzing current evaluation reports (Section 4), and analyzing evaluation frameworks (Section 5), we make progress toward a more contextual and critical understanding of AI audits.
@@ 3 What information do audit cards need for rigorous reporting?
Methodology: We identify key components of audit cards through an initial literature review of work on AI audits and evaluations, transparency, technical evaluation design, and sociotechnical approaches. [^3] We also refined our audit card template based on structured interviews with expert stakeholders. See Section 6. We selected 28 total papers and manually annotated them, scoring the extent to which the paper recommends that aspect of reporting: is it a major argument of the paper (2), a minor argument of the paper (1), or not mentioned in the paper (0). See Appendix B for further details on the methodology. In Table 1, we summarize the perspectives from all 28 papers on the three principles and six features. It’s aim is to be a comprehensive and exhaustive overview of audit features based on the relevant literature. Next, we expand on these principles and features. See Figure 1 for an overview of an audit card and Appendix C for an actionable audit card checklist that discusses the relative importance of principles and features for different stakeholders and auditing types.
Principles Features
Paper Justif. Assum. Limit. Who What How Access Integrity Review
Anderljung et al.2023 1 2 2 2 2 2 2 2 2
Apollo Research2024 2 2 0 0 2 2 0 0 0
Barnett and Thiergart2024a 2 2 0 1 0 0 0 0 0
Barnett and Thiergart2024b 1 2 2 0 2 2 2 2 0
Birhane et al.2024 2 2 2 2 2 0 1 2 2
Bucknall and Trager2023 0 2 2 0 2 0 2 2 2
Burnell et al.2023 2 2 0 0 1 2 0 0 2
Casper et al.2024 1 2 0 1 0 0 2 0 1
Chang et al.2023 0 2 0 2 0 2 0 0 0
Costanza-Chock, Raji, and Buolamwini2022 2 2 0 2 1 0 2 1 2
Dobbe, Krendl Gilbert, and Mintz2021 2 2 2 2 1 2 0 0 2
Dow et al.2024 0 2 0 1 2 2 0 0 0
Eriksson et al.2025 2 2 1 1 2 2 0 1 0
Gallifant et al.2025 2 2 2 2 2 2 1 2 2
Gebru et al.2021 1 2 2 2 1 2 2 2 2
Hendrycks and Woodside2024 0 0 1 0 0 2 0 0 1
Kolt et al.2024 1 1 0 2 1 0 0 2 2
Liang et al.2023 1 2 0 0 2 0 0 0 0
METRb 0 0 0 2 2 2 0 0 2
Mökander et al.2024 0 2 2 1 2 1 2 2 2
Mukobi2024 0 2 0 2 2 2 2 2 1
Ojewale et al.2024 1 2 0 2 1 1 2 1 1
Raji et al.2020 0 2 2 2 2 2 1 0 1
Reuel et al.2024b 0 2 2 2 2 2 0 0 2
Selbst et al.2019 0 2 2 2 0 2 0 0 0
Shevlane et al.2023 0 2 2 2 2 2 2 2 2
Weidinger et al.2025 2 2 2 0 2 2 0 0 1
Zhan et al.2024 2 2 2 1 2 2 0 0 2
@@
Table 1: What contextual details do prior works say are key for audit reporting? We divide reporting into three overarching principles (justifications, assumptions, limitations) and six key features (who, what, how, access, integrity, review). We score each paper’s extent of recommendation of each these as 2 (major argument), 1 (minor argument), and 0 (no mention). Refer to Table 4 for our complete, more granular analysis of papers.
Overarching principles for transparent reporting
@@
The following three overarching principles behind auditing are key for transparency and methodological clarity. These apply across the entire auditing process and the six specific features. We design audit cards to make it easier to adopt these principles throughout the auditing process. This, in return, enables accurate interpretation, facilitates constructive critique, and builds trust in the evaluation process.
@@
Justifications as explicit arguments to support methodological choices, metric selection, and interpretative frameworks. Explicit reporting on justifications enables process transparency. This justification can be through elements of a rigorous scientific process, such as explaining proxies and experimental coverage (Apollo Research 2024), or through standardized auditing procedures, such as detailed standards identification (Birhane et al. 2024; Costanza-Chock, Raji, and Buolamwini 2022).
@@
Assumptions to articulate the premises underlying the evaluation design and analysis. [^4] This differs from justifications insofar that they answer “Why did we choose this approach?” whereas assumptions answer “What are we taking as given?” Reporting on assumptions clarifies the relationship (or lack thereof) between tests and real-world outcomes, presumptions about what constitutes “good” performance, and details about the threat models considered (Reuel et al. 2024b; Barnett and Thiergart 2024a, b; Raji et al. 2020).
@@
Limitations to explicitly acknowledge the constraints that affect the validity, reliability, or generalizability of findings. Reporting limitations clarifies constraints in data sampling, potential artifacts in the evaluation process, and boundaries of what auditors can reasonably claim based on the evaluation evidence (Mökander et al. 2024; Barnett and Thiergart 2024b; Reuel et al. 2024b).
Specific features for contextual information
@@
Aside from overarching principles, reporting on more specific contextual features of audits helps stakeholders assess trustworthiness. We compile recommendations from past literature into six features for reporting evaluation context in audit cards. For a more granular analysis, see Appendix C, where we further divide these six features into constituent aspects.
@@
  • Who are the auditors: Reporting details of auditor identity facilitates trust and clarifies what potential biases might affect their work. It includes reporting the expertise (relevant domain knowledge, accreditations) (Anderljung et al. 2023; METR b; Reuel et al. 2024b) and background (track record of conducting evaluations, positionality) of auditors (Costanza-Chock, Raji, and Buolamwini 2022; Gebru et al. 2021). Given worries about the privacy of auditors, information about the background can be anonymized and aggregated.
@@
  • What is evaluated: Reporting on evaluation goals enables clear interpretation of findings. This includes the scope (e.g., model vs. scaffolding) and goal (context of use) (Raji et al. 2020; Birhane et al. 2024; Shevlane et al. 2023), type of evaluation (e.g., capability vs. propensity), as the scope of the evaluation influences what the relevant level of access is (METR 2024; Liang et al. 2023), as well as reporting on the evaluation obsolescence criteria (Joaquin et al. 2025). [^5] Examples of obsolescence criteria could include substantial amounts of model fine-tuning, changes in system components, or changes in deployment context. This helps stakeholders understand the shelf life of the findings. It also includes reporting whether the evaluation will be repeated, and if so, at what time or after which conditions are met (Barnett and Thiergart 2024b; Anderljung et al. 2023).
@@
  • How is it evaluated: Explaining the underlying procedures for the evaluation, such as descriptions of the evaluation setup with justifications of how a capability translates to evaluation goal (Apollo Research 2024; Raji et al. 2020; Reuel et al. 2024b) and provide interpretation of scores (Dobbe, Krendl Gilbert, and Mintz 2021; Selbst et al. 2019). The specific relevant aspects to report for this feature may change depending on the underlying evaluation. We therefore refer to the appropriate literature for the underlying evaluation for a complete list of relevant aspects to report and details on how to report on them, e.g. for benchmarks (Reuel et al. 2024b).
@@
  • Access and resources: Report the constraints under which the evaluation was conducted to allow readers to critically assess the audit’s limitations. This includes, what access auditors have to a system (black/white box, with/without safeguards) and which resources available are available (compute, time, funding) as this can have a large impact on the rigor of evaluations (Barnett and Thiergart 2024b; Ojewale et al. 2024; Casper et al. 2024).
@@
  • Integrity: Accountability for the auditing process establishes trust in auditors and rules out potential conflicts of interest. This includes reporting transparently about the selection process (how auditors were chosen) and compensation (funding sources, incentive structures). These dynamics can significantly affect the evaluation process and are therefore crucial context (Birhane et al. 2024; Costanza-Chock, Raji, and Buolamwini 2022; Gebru et al. 2021).
@@
  • Review and communication: Reports on if and what independent reviews of audit methods facilitate accountability [^6] We discuss issues of intellectual property and non-disclosure agreements (NDAs) standing in the way of internal review in Section 6 and Section 7. This review can be performed by experts at the organization by cross-checking the final report, insofar as they were not involved in the evaluation beforehand. (Reuel et al. 2024b; Anderljung et al. 2023). This further includes contact information and avenues for corrections, executive summaries (covering the evaluation process and result for non-technical audiences), and clear definitions of key terms (Gebru et al. 2021; Bucknall and Trager 2023; METR 2024; Reuel et al. 2024b).
@@
Some of the above features may be considered confidential in certain contexts, with academic researchers potentially sharing different details than commercial entities (Mitchell et al. 2019). Recommendations about how and with whom audit cards should be shared are beyond the scope of this report. However, redactions offer a simple, well-precedented approach for protecting sensitive information (Boone, Floros, and Johnson 2015).